OpenStudy (anonymous):

I am completely lost... 4b - 1/7 = 5/7 How do I solve this?

OpenStudy (anonymous):

4b= 5/7+1/7 b= (5/7+1/7)/4

OpenStudy (anonymous):

._. but I dunno how to get b isolated.

OpenStudy (anonymous):

I have just isolated be for you

OpenStudy (anonymous):

Can you tell me how? lol I suck with fractions.

OpenStudy (anonymous):

you will need to move the like terms to the same side of the equation.in this case the fractions to one side. moving the negative acroos the equal sign will make it positve

OpenStudy (anonymous):

Like terms being 1/7 and 5/7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so moving the -1/7 across makes it positive so you add 1/7 to the other fraction

OpenStudy (anonymous):

and to isolate b since b is multiplying 4 you will move the to the other side of the equation. It will go to divide the sum of the fraction on the right!

OpenStudy (anonymous):

i meant move the 4 to the other side

OpenStudy (anonymous):

OH so 6/7 divided by 4?

OpenStudy (anonymous):

I completely forget if one is too add denominatiors.

OpenStudy (anonymous):

Anyway, I get it more now, thank you :)

OpenStudy (anonymous):

If the denominators are the same, you add the numerators

OpenStudy (anonymous):

Ah okay.

OpenStudy (anonymous):

I get it now with this format, tyvm :D I understand what you mean with b = (5/7 + 1/7)/4